A Prius moving at a speed of 20 m/s begins to slow down at a constant rate of 3.0 m/s each second. Find how far it goes before stopping. (Remember that when the Prius stops, its final velocity =0)
Answer #1

Using the equation of motion :

v^2= u^2 + 2as

v = 0

u = 20 m/s

a = -3 m/s2

So, 0^2 = 20^2 + 2*(-3)*s

So, s = 66.7 m <-------- answer